OpenStudy (tylerd):

show y''=-y^-3 on the circle x^2+y^2=1

OpenStudy (tylerd):

\[x^2+y^2=1\] \[y''=-y^{-3}\]

OpenStudy (tylerd):

need to prove

OpenStudy (freckles):

have you tried to differentiate x^2+y^2=1?

OpenStudy (tylerd):

yes

OpenStudy (freckles):

well you first need to find y' then find y''

OpenStudy (freckles):

let me know if you come back

OpenStudy (freckles):

and need any help

OpenStudy (tylerd):

2x+2yy'=0 2x=-2yy' x=-yy'

OpenStudy (freckles):

so y'=x/-y

OpenStudy (freckles):

so find y'' by differentiating both sides

OpenStudy (freckles):

use quotient rule for the (-x/y) part

OpenStudy (tylerd):

i did

OpenStudy (tylerd):

but didnt get the answer

OpenStudy (freckles):

what did you get

OpenStudy (tylerd):

can you show me your steps?

OpenStudy (freckles):

anyways if you found the correct y'' then you will replace y' with -x/y and then get rid of the compound fraction then use that x^2+y^2=1

OpenStudy (freckles):

then you will have your -1/y^3=y''`

OpenStudy (freckles):

did you want to show me what you got for y''

OpenStudy (freckles):

because as I was saying you have to manipulate what you have for y'' to get the -1/y^3

OpenStudy (tylerd):

im still not getting it

OpenStudy (tylerd):

@SithsAndGiggles

OpenStudy (anonymous):

@freckles has everything you need. Differentiating the circle equation, you have \[\frac{d}{dx}[x^2+y^2=1]~~\implies~~2x+2yy'=0~~\implies~~y'=-\frac{x}{y}\] Differentiating again, you have (via the quotient rule) \[y''=-\frac{y-xy'}{y^2}\] Substitute \(y'=-\dfrac{x}{y}\), and you have \[y''=-\frac{y+x\dfrac{x}{y}}{y^2}~~\iff~~y''=-\frac{y+\dfrac{x^2}{y}}{y^2}\] Multiply the RHS by \(\dfrac{y}{y}\), and you have \[y''=-\frac{y^2+x^2}{y^3}\] You know that \(x^2+y^2=1\), so \[y''=-\frac{1}{y^3}=-y^{-3}\] Notice that I've done nothing different from above.